OpenStudy (anonymous):

Meg has can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. Part B: how many liters of orange juice are present in the can? Show your work

OpenStudy (danjs):

1 Liter of water is 20% of the total volume 1 L = 0.20*(Total)

OpenStudy (anonymous):

So is that for part a or part b @DanJS

OpenStudy (danjs):

um.

OpenStudy (danjs):

Yeah that is part A, they want an equation to determine the total volume 1 = 0.20*V where V is the total volume

OpenStudy (danjs):

so V = 1/0.2 = 5 The total is 5 Liters, 1 Liter is Water, so, 4 Liters is Juice

OpenStudy (anonymous):

Thank you @DanJS